Review Part A You go to an amusement park with your friend Betty. who wants to ride the 60-n-diameter Ferris wheel. She starts the ride at the lowest point of a wheel that as you face it, rotates What is the magnitude of her displacement vector when the wheel has rotated by an angle of 80? Express your answer with the appropriate units. ArValue Units Submit REusL Ansver Part B What is the direction of her displacement vector when the wheel has rotated by an angle of 80, measured counterclockwise from the horizontal? Submit Request Answer
Answer #1

Here, let's take horizontal as the x axis

and vertical as the y axis

for the initial position

ri = -30 j m

for the final speed ,

rf = 30 * (sin(80) i - j *cos(80)) = 29.5 i - 5.2 j m

for the displacement vector

r = rf - ri

r = 29.5 i - 5.2 j - (-30 )j

r = 29.5 i + 24.8 j

magnitude of displacement vector = sqrt(29.5^2 + 24.8^2)

magnitude of displacement vector = 38.34 m

part B)

direction of her displacement = arctan(24.8/29.5)

direction of her displacement = 40 degree